A vehicle travels 280 miles in 5 hours. At this rate, how many miles will the vehicle travel in 8 hours?
Answer by oberobic(2304)

Start with the basic distance equation: d = r*t
.
d = 280 miles
t = 5 hr
r = 280/5 = 56 miles/hr
.
8*56 = 448 miles in 8 hrs at 56 mph
.
Done.